# Forests of Uncertaint(r)ees: Using tree-based ensembles to estimate probability distributions of future conflict

### v1

@@ 3 Methods
As the model was submitted to the VIEWS prediction challenge (Hegre et al., 2025), we generated predictions for each month in six yearly test windows (2018-2023) and (at the time) true future predictions for July 2024-June 2025.333The true future predictions are tracked and evaluated at https://viewsforecasting.org/research/prediction-challenge-2023/leaderboard/.
@@ 3.1 Data and Modeling Setup
With our main focus on modeling strategies, we rely on the data provided by the VIEWS (Violence & Impacts Early-Warning System) team in the context of the challenge and use all features provided [^4] These include information on conflict history, geography, natural resources, population size, climate and vegetation, and to some extent the economy. A codebook is available with the data provided in the context of the VIEWS challenge (Hegre et al., 2023). in our models. The data covers all PRIO-GRID cells in Africa and the Middle East, a subset with N=13110 cells of the global 0.5° x 0.5° grid (Tollefsen et al., 2012), and are available monthly starting in 1990. As outlined in the competition call, the target is the number of fatalities from state-based armed conflict events (Hegre et al., 2023), as recorded by the Uppsala Conflict Data Program (UCDP) (Sundberg and Melander, 2013; Davies et al., 2025). The target is highly zero-inflated, with less than 0.4% non-zero values in the average month in our training data.
@@
To generate predictions for the whole next year from the available training data, we chose to train separate models for each of the timesteps to predict (t+3, … t+14) and combine the resulting outputs to a full year of predictions. The data cutoff for each test window is 3 months before the start, i.e. October 2017 for 2018 predictions. [^5] For each of the six yearly test windows, the training data is artificially limited to October of the previous year. This is done to simulate the availability gap present in the data for the true future predictions, where training data was available until April 2024 to predict for the timeframe July 2024 to June 2025. Therefore, predictions need to be generated for the timesteps t+3, …, t+14. The training data is further limited by the time period we want to predict into the future with this approach. For example, the model generating the predictions for December 2018 (t+14) based on data up to October 2017 (t) can only be trained on data up to August 2016 (t-14), as this is the last month where there is sufficient future information to label the target.
@@
We designed a modular, model-agnostic modeling pipeline in Python, which performs the tuning, training and predicting automatically for the given prediction problem. We include multiple machine learning algorithms, selecting the algorithm which achieves the best performance during cross-validation for each timestep individually. This allows us to integrate and test different machine learning algorithms with minimal effort and means our approach can be easily reused for different prediction problems, with the code available as part of our replication material.
@@
We perform hyperparameter tuning for each timestep once based on the data up to October 2017 (N=4,378,740), before the first test window. Our tuning procedure is based on time series cross-validation with a 5-year sliding window through the training data. We employ the hyperopt package (Bergstra et al., 2013), which implements a Bayesian search approach with the Tree-structured Parzen Estimator (TPE) algorithm (Bergstra et al., 2011), to identify the best hyperparameters for each prediction timestep model. The cross-validation and scoring within the tuning procedure is implemented in the scikit-learn framework (Pedregosa et al., 2011). After tuning, we fit the model on all available training data for a given test window before generating our predictions from the last available observations 3 months before the start of the prediction window.
@@
Our tuning metric depends on the modeling step in the hurdle model, with further details provided below. For each chosen base metric, our tuning metric is defined as the mean performance across the test folds during cross-validation. As we observed a strong tendency to overfit for some parameter combinations, we combine this with a penalty term on deviations between mean train and test fold performance, in order to favor generalizability of the models. [^6] The formula we use is $M_{tune}=\overline{M}_{test}-0.5|\overline{M}_{train}-\overline{M}_{test}|$, where higher scores for all $M$ are better, and $\overline{M}$ represents the mean performance across all train respectively test folds in the cross-validation. The tuning metric $M_{tune}$ is then maximized by the TPE tuning algorithm.
@@ 3.2 Hurdle Approach
Our overarching modeling approach is a variation on the principle of hurdle models. Hurdle models (Mullahy, 1986) are a combination of two modeling steps: a first step consisting of a classifier determining whether the hurdle is reached, trained on all available training data, and a second step consisting of a regressor determining the predicted value, trained on only the subsection of the training data where the target has reached the hurdle. We perform tuning and prediction for each of these two steps separately to allow for custom combinations of global and local predictions (see below).
@@ 3.2.1 Classifier
Our classifiers are trained on a binary variable indicating whether or not any fatalities occurred in a given pgm. We include two different probabilistic classifiers in our modeling pipeline, Random Forests (Breiman, 2001) and eXtreme Gradient Boosting (Chen and Guestrin, 2016) models. [^7] We also tested logit models as a “simple” alternative approach, which performed worse by a factor of 2-3 on average.
  • Random Forests (RF) estimate probability by aggregating the results of a multitude of decision trees. Each tree in the forest is built from a different sample of data, using a technique called bootstrap aggregating, or bagging. Additionally, Random Forests employ random feature selection, where each split in a tree considers only a random subset of features.
  • eXtreme Gradient Boosting (XGB), employs a regularized learning objective that balances model complexity and predictive accuracy. The system utilizes gradient tree boosting, where the model is trained in an additive manner, incrementally improving the predictions by minimizing a loss function using second-order gradient statistics. Models are built sequentially, correcting the errors of previous models.
Our tuning metric of choice for classification is the average precision score, which summarizes the precision-recall-curve. This metric is well suited for zero-inflated classification tasks as it does not take into account whether zeroes are predicted correctly, which we argue is of little interest given the relative scarcity of violence (also see Saito and Rehmsmeier, 2015). Tuning performance for both model types is fairly similar, with RF performing slightly better in all global models, while XGB is favored for most local models.
@@ 3.2.2 Regressor
To estimate uncertainty around predictions, we rely on regressors designed to output distributions directly rather than trying to estimate distributions around point predictions ex post. We include three different tree-based distributional regressors in our modeling pipeline, Quantile Regression Forests (Meinshausen, 2006), Distributional Random Forests (Cevid et al., 2022) and Natural Gradient Boosting for probabilistic regression (Duan et al., 2019).
@@
  • Natural Gradient Boosting (NGB) for probabilistic regression extends gradient boosting to the estimation of probability distributions. This involves boosting the parameters of a specified parametric distribution using a natural gradient, which corrects the training dynamics for more stable and efficient learning. The algorithm integrates three modular components: a simple base learner, a parametric probability distribution, and a proper scoring rule. The natural gradient is employed to optimize the parameters of the conditional distribution, ensuring that the updates are invariant to reparameterization and efficiently exploit the curvature of the score in distributional space. We use decision trees as the base learner, log-normal probability distributions and the CRPS as the scoring rule.
Following the principle of hurdle models, we train our regression models only on pgms with non-zero targets while still generating predictions for all pgms. As our tuning metric, we use the competition’s main metric, the CRPS (Hegre et al., 2025). In line with the maximum number of samples allowed by the prediction challenge, we set our regression models to output 1000 samples of the predicted distribution. This results in a wider range of possible values, ensuring the inclusion of low-probability outcomes. NGB performed best during tuning, being chosen in 75% of global models and 80% of local models, with the other two regressors only chosen occasionally.
@@ 3.2.3 Quasi-Hurdle Ensemble
Hurdle model point predictions are usually generated via a simple multiplication of the output of the classifier and the predicted value of the regressor. Given that we work with N=1000 samples drawn from the predictive distribution instead, a multiplication of the classification probability with each of the samples would result in non-integer predictions [^8] The only exception here is NGB which does also predict zeroes. For consistency with the other algorithms we replace all zeros with ones in the NGB predictions., which is not in line with the nature of fatality counts. At the same time, our tree-based regressors trained only on non-zero targets never produce zero predictions and a multiplication would therefore likely overestimate the probability of violence occurring. While both issues could be partially addressed with rounding, we opt to instead interpret the classification probability as the percentage of the ensemble sample taken from the non-zero predictions via a random draw, with the remaining share of the 1000 samples filled with zero values. In testing, this also performed better than the multiplicative approach. [^9] We also tested selecting either all-zero samples or the full non-zero sample based on the predicted probability from the classifier and a threshold, which performed significantly worse.
@@ 3.3 Local Models
In addition to the “global” approach described above, we explore the integration of multiple “local” models encompassing smaller geographic regions. This has two potential benefits: first, it reduces the coverage requirements and thus increases the pool of available datasets for a prediction problem. Second, it accounts for potential biases across different regions, either due to differences in conflict patterns, or due to differences in data generation. Focussing on potential differences in conflict patterns, we create custom contiguous geographic clusters based on the spatial distribution of grid cells with any recorded fatalities in the training data (1990-2017) using the HDBSCAN clustering algorithm (Campello et al., 2013) [^10] We tested two additional versions of generating clusters: One with clusters created via an alternative clustering algorithm, DBSCAN, and a similar manual tuning of clustering parameters, and one with clusters corresponding to the United Nations Statistics Division sub-regions more aligned to the data coverage problem, with grid cells assigned to countries based on a majority rule. Both performed only slightly worse in testing. While we did not include them in our final model run and evaluation, corresponding models can still be produced with our replication code.. We manually tested parameter combinations for the clustering algorithm until inspections of the results yielded groupings, which plausibly corresponded to visually discernible patterns, resulting in eleven clusters. To ensure sufficient non-zero training data in each cluster for the hurdle regression models, we further reduce this down to six clusters by iterating over the clusters and combining smaller clusters with their nearest neighbors based on centroid distance of polygons drawn around each cluster, requiring a minimum of 1000 pgms with non-zero fatalities.
To subsequently assign any cells not containing conflict, we first draw new polygons around the combined grid cells of each cluster. Cells remaining outside these polygons were assigned to the cluster with the nearest boundary to the cell center. The procedure and resulting clusters are visualized in Figure 1. We subsequently train separate “local” models for each of the clusters following the same procedure as with the “global” models, described above. Each grid cell is therefore assigned not only to exactly one cluster but also to one corresponding local set of models. Combining the predictions from all local models yields predictions for the whole geographical area of interest.
@@ 3.3.1 Global-Local Ensemble
We create two spatial ensembles in addition to the original global hurdle model: a local-only hurdle model following the same approach, simply concatenating the predictions from all local models along their spatial IDs, and a global-local ensemble. The latter is constructed by selectively picking and choosing both classification and regression components from either the global or the local model, creating an ensemble of global and local model outputs. We select the components by comparing the combined performance of the hurdle ensemble for each of the four possible combinations of classification and regression predictions on a cluster-by-cluster basis, selecting the global-local combination for each cluster which performed best across the three years prior to a given prediction window. [^11] We also tested this using only one or two years of prior data, which resulted in worse performance of the combined prediction. This is likely connected to a fairly high volatility in performance across years, with prior performance not correlated enough to future performance. An increase to five years of prior data as the selection basis did not lead to meaningful improvements. Finally, the resulting predictions for each cluster are spatially concatenated analogous to the local-only model.
Figure 1: Visualization of the creation process of the clusters for local models. a) clusters created by manually tuned HDBSCAN algorithm with corresponding polygons; b) clusters after merging of smaller clusters with updated polygons; c) final clusters with grid-cells without violence assigned. Grid cells shown in a) and b) are those experiencing any violence in the training data (1990-2017). Black grid cells displayed in a) and b) are not assigned to any clusters by HDBSCAN initially.
4 Results
To illustrate what kind of intervals our model predicts, Figure 2 compares observed fatalities during the first test window with our predictions for four example grid cells representing typical prediction patterns: A high-intensity case (a) and a low-intensity case (b) we capture fairly well, and two “misses”: a “false negative” case (c) where our predictions showed only a minuscule chance for violence but in fact a significant number of fatalities occurred, and a “false positive” case (d) where our predictions saw a small but noticeable chance for violence usually seen in low intensity cases, but no violence occurred.
As is immediately apparent, our predictions are heavily right-skewed with long tails and a number of zeroes, indicating mostly low probabilities in the classification step. Only small prediction intervals for the highest-intensity predictions have lower bounds above zero (roughly 25% and smaller, Figure 2a). Notably, these characteristics are also found in the distribution of observed fatalities across pgms. The predictions capture general conflict intensity well in most cases, with the distributional approach treating the spikes often observed in real-world data simply as lower likelihood events, but within the range of possible values. Typical misses include “false negative” cases, where significant violence occurred despite our sample containing more than 99% zeroes, such as in Cameroon in the vicinity of the border to Nigeria (Figure 2c). However, as also illustrated, our models always assume a small base probability of violence with the samples never being all-zero, which is the result of our hurdle combination approach and a non-zero minimum probability from the classification step. Lastly, we see a number of “false positives” (Figure 2d), where a sizable share of our samples is non-zero but this predicted risk does not materialize. These predictions generally are part of a spatial cluster of higher-likelihood predictions with fatalities observed only in some of its grid cells. The distributional approach thus highlights elevated risk in cells near active conflict and predictions are similar to those where low-intensity violence did occur (Figure 2b), without being able to reliably identify exact locations.
@@ 4 Results
Figure 2: Prediction intervals for example grid cells based on 2018 predictions from the global model, compared to the observed number of fatalities and a maximum a posteriori (MAP) estimate for the predictive distributions. MAP estimates are calculated as the maximum of the probability density function estimated via Gaussian kernel density estimation. 25%, 50%, 75%, 90%, and 95% prediction intervals are only drawn if upper boundaries are not zero. Figures 2b) and 2c) also include the extreme boundaries (min/max) of the corresponding samples.

### v2

@@ 3 Methods
As the models were submitted to the VIEWS prediction challenge (Hegre et al., 2025), we generated predictions for each month in six yearly test windows (2018-2023) and (at the time) true future predictions for July 2024-June 2025.333The true future predictions are tracked and evaluated at https://viewsforecasting.org/leaderboard/ (VIEWS, 2026).
@@ 3.1 Data and Modeling Setup
With our main focus on modeling strategies, we rely on the data provided by the VIEWS team in the context of the prediction challenge and use the full feature set [^4] These include information on conflict history, geography, natural resources, population size, climate and vegetation, and to some extent the economy. A codebook is available with the data provided in the context of the VIEWS challenge (Hegre et al., 2023). in our models. The data covers all PRIO-GRID (Tollefsen et al., 2012) cells in Africa and the Middle East, a subset with $N=13,110$ cells of the global $0.5^{\circ}\times 0.5^{\circ}$ grid, and are available monthly starting in 1990. As outlined in the prediction challenge call, the target is the number of fatalities from state-based armed conflict events (Hegre et al., 2023), as recorded by the Uppsala Conflict Data Program (UCDP) (Sundberg and Melander, 2013; Davies et al., 2025). The target is highly zero-inflated, with less than 0.4% non-zero values in the average month in our training data.
@@
To generate predictions for the whole next year from the available training data, we chose to train separate models for each of the timesteps to predict (t+3,,t+14t+3,...,t+14) and combine the resulting outputs to a full year of predictions. The data cutoff for each test window is three months before the start, i.e. October 2017 for 2018 predictions. [^5] This simulates the availability gap present in the data for the true future predictions, where training data was available until April 2024 to predict for the timeframe July 2024 to June 2025. Therefore, predictions need to be generated for the timesteps t+3,,t+14t+3,...,t+14. The training data is further limited by the time period we want to predict into the future with this approach. For example, the model generating the predictions for December 2018 ($t+14$) based on data up to October 2017 ($t$) can only be trained on data up to August 2016 (t14t-14), as this is the last month where there is sufficient future information to label the target.
@@
We designed a modular, model-agnostic modeling pipeline to perform the tuning, training, and predicting automatically for the given prediction problem. We include multiple machine learning algorithms, selecting the algorithm which achieves the best performance during cross-validation for each timestep individually. This allows us to integrate and test different machine learning algorithms with minimal effort and means our approach can be easily reused for different prediction problems, with the code available as part of our replication material.
@@
We perform hyperparameter tuning for each timestep once based on the data up to October 2017 ($N=4,378,740$), before the first test window. Our tuning procedure is based on time series cross-validation(Petropoulos et al., 2022) with a 5-year sliding window through the training data. We employ a Bayesian search approach with the Tree-structured Parzen Estimator (TPE) algorithm using hyperopt (Bergstra et al., 2011, 2013), to efficiently estimate the best hyperparameters for each prediction timestep model. The cross-validation and scoring within the tuning procedure is implemented in the scikit-learn framework (Pedregosa et al., 2011). After tuning, we fit the model on all available training data for a given test window before generating our predictions from the last available observations three months before the start of the prediction window.
@@
Our tuning metric depends on the modeling step in the hurdle model, with further details provided below. For each chosen base metric, our tuning metric is defined as the mean performance across the test folds during cross-validation. As we observed a strong tendency to overfit for some parameter combinations, we combine this with a penalty term on deviations between mean train and test fold performance, in order to favor generalizability of the models. [^6] The formula we use is $M_{tune}=\overline{M}_{test}-0.5|\overline{M}_{train}-\overline{M}_{test}|$, where higher scores for all $M$ are better, and $\overline{M}$ represents the mean performance across all train and test folds in the cross-validation, respectively. The tuning metric $M_{tune}$ is then maximized by the tuning algorithm.
@@ 3.2 Hurdle Approach
Our overarching modeling approach is a variation on the principle of hurdle models. Hurdle models (Mullahy, 1986) are a combination of two modeling steps: a first step consisting of a classifier determining whether the hurdle is reached, trained on all available training data, and a second step consisting of a regressor determining the predicted value, trained on only the subsection of the training data where the target is above the hurdle. We perform tuning and prediction for each of these two steps separately to allow for custom combinations of global and local predictions (see 3.2.3).
@@ 3.2.1 Classifier
Our classifiers are trained on a binary variable indicating whether or not any fatalities occurred in a given pgm. We include two different probabilistic classifiers in our modeling pipeline: Random Forests (Breiman, 2001) and eXtreme Gradient Boosting (Chen and Guestrin, 2016) models. [^7] We also tested logit models as a “simple” alternative approach, which performed worse by a factor of 2-3 on average.
  • Random Forests (RF) aggregate the results of a multitude of decision trees. Each tree in the forest is built from a different sample of data, using a technique called bootstrap aggregating, or bagging. Additionally, Random Forests employ random feature selection, where each split in a tree considers only a random subset of features.
  • eXtreme Gradient Boosting (XGB) employs a regularized learning objective that balances model complexity and predictive accuracy. The system utilizes gradient tree boosting, where the model is trained in an additive manner, incrementally improving the predictions by minimizing a loss function using second-order gradient statistics. Models are built sequentially, correcting the errors of previous models.
Our tuning metric of choice for classification is the area under the precision-recall-curve. This metric is well suited for zero-inflated classification tasks, as it does not take into account whether zeroes are predicted correctly – which we argue is of little interest given the relative scarcity of violence (also see Saito and Rehmsmeier, 2015). Tuning performance for both model types is fairly similar, with RF performing slightly better in all global models, while XGB is favored for most local models.
@@ 3.2.2 Regressor
To estimate uncertainty around predictions, we rely on regressors designed to output distributions directly rather than trying to estimate distributions around point predictions ex post. We include three different tree-based distributional regressors in our modeling pipeline: Quantile Regression Forests (Meinshausen, 2006), Distributional Random Forests (Cevid et al., 2022) and Natural Gradient Boosting for probabilistic regression (Duan et al., 2019).
@@
  • Natural Gradient Boosting (NGB) for probabilistic regression extends gradient boosting to the estimation of probability distributions. This involves boosting the parameters of a specified parametric distribution using a natural gradient, which corrects the training dynamics for more stable and efficient learning. The algorithm integrates three modular components: a simple base learner, a parametric probability distribution, and a proper scoring rule. The natural gradient is employed to optimize the parameters of the conditional distribution, ensuring that the updates are invariant to reparameterization and efficiently exploit the curvature of the score in distributional space. We use decision trees as the base learner, log-normal probability distributions and the Continuous Ranked Probability Scores (CRPS) as the scoring rule.
Following the principle of hurdle models, we train our regression models only on pgms with non-zero targets while still generating predictions for all pgms. As our tuning metric, we use the challenge’s main metric, the CRPS (see equation 1, Hegre et al., 2025). In line with the maximum sample size allowed by the prediction challenge, we set our regression models to output a forecast sample with $N_{sample}=1,000$, drawn from the predicted distribution. This results in a wider range of possible values, ensuring the inclusion of low-probability outcomes. NGB performed best during tuning, being chosen in 75% of global models and 80% of local models, with the other two regressors only chosen occasionally.
Figure 1: Visualization of the process to create the geographic regions for local models at three steps: a) clusters created by the manually tuned HDBSCAN algorithm with corresponding polygons; b) updated clusters after combining smaller clusters with corresponding polygons; c) final contiguous regions with grid-cells without violence integrated. Grid cells shown in a and b are those experiencing any violence in the training data (1990-2017). Black grid cells displayed in a and b are not assigned to any clusters by HDBSCAN initially.
@@ 3.2.3 Quasi-Hurdle Ensemble
Hurdle model point predictions are usually generated via a simple multiplication of the output of the classifier and the predicted value of the regressor. Given the samples drawn from the predictive distribution, a multiplication of the classification probability with each draw in the samples would result in non-integer predictions, which is not in line with the nature of fatality counts. At the same time, our tree-based regressors trained only on non-zero targets never produce zero predictions [^8] The only exception here is NGB which does also predict zeroes. For consistency with the other algorithms we replace all zeros with one as the lowest possible non-zero value in the NGB predictions., a multiplication would therefore likely overestimate the probability of violence occurring. While both issues could be partially addressed with rounding, we opt to instead treat the classification probability as the percentage of the ensemble sample taken from the non-zero predictions via a random draw, with the remaining share of the $1,000$ draws filled with zero values. In testing, this also performed better than the multiplicative approach. [^9] We also tested a more ”traditional” hurdle combination, selecting either all-zero samples or the full non-zero sample based on the classification probability and a threshold, which performed significantly worse.
@@ 3.3 Local Models
In addition to the “global” approach described above, we explore the integration of multiple “local” models encompassing smaller geographic regions. This has two potential benefits: first, it reduces the spatial coverage requirements and thus increases the pool of available datasets for a prediction problem. Second, it accounts for potential systematic differences across regions, either due to differences in conflict patterns or due to differences in data generation. Focusing on potential differences in conflict patterns, we create custom contiguous geographic regions based on the spatial distribution of grid cells with any recorded fatalities in the training data (1990-2017) using the HDBSCAN clustering algorithm (Campello et al., 2013). [^10] We tested two additional versions of generating clusters: One with clusters created via an alternative clustering algorithm, DBSCAN, and a similar manual tuning of clustering parameters, and one with clusters corresponding to the United Nations Statistics Division sub-regions more aligned to the data coverage problem, with grid cells assigned to countries based on a majority rule. Both performed only slightly worse in testing. While we did not include them in our final model run and evaluation, corresponding models can still be produced with our replication code. We manually tested parameter combinations for the clustering algorithm until inspections of the results yielded groupings, which plausibly corresponded to visually discernible patterns, resulting in eleven clusters. To ensure sufficient non-zero training data in each cluster for the hurdle regression models, we further reduce this down to six clusters by iterating over the clusters and combining smaller clusters with their nearest neighbors based on centroid distance of polygons drawn around each cluster, requiring a minimum of $1,000$ pgms with non-zero fatalities.
To subsequently create contiguous regions including any cells not experiencing conflict, we first draw new polygons around the combined grid cells of each cluster. Cells remaining outside these polygons were assigned to the cluster with the nearest boundary to the cell center. The procedure and resulting clusters are visualized in Figure 1. We subsequently train separate local models for each of the clusters, following the same procedure as with the global models described above. Each grid cell is therefore assigned not only to exactly one cluster but also to one corresponding local set of models. Combining the predictions from all local models yields predictions for the whole geographical area of interest.
@@ 3.3.1 Global-Local Ensemble
We create two spatial ensembles in addition to the original global hurdle model: a local-only hurdle model following the same approach, simply concatenating the predictions from all local models along their spatial IDs, and a global-local ensemble. The latter is constructed by selectively picking and choosing both classification and regression components from either the global or the local model, creating an ensemble of global and local model outputs. We select the components by comparing the combined performance of the hurdle ensemble for each of the four possible combinations of classification and regression predictions on a cluster-by-cluster basis, selecting the global-local combination for each cluster which performed best across the three years prior to a given prediction window. [^11] Repeating this with only one or two years of prior data resulted in worse performance of the combined prediction. This is likely connected to a fairly high volatility in performance across years, with prior performance not correlated enough to future performance. An increase to five years of prior data as the selection basis did not lead to meaningful improvements. Finally, the resulting predictions for each cluster are spatially concatenated analogous to the local-only model.
@@
Figure 2: Predictive distributions and their development over time for example grid cells in the global model’s 2018 predictions, compared to the observed number of fatalities and a maximum a posteriori (MAP) estimate for the distributions. Figures a and c show the distributions for single pgms, estimated via Gaussian kernel density estimation from the forecast samples, with X-axes limited to below the 98% quantile. Figures b, d-f show the times series for the whole test window, visualized through selected quantile-based prediction intervals. MAP estimates are calculated as the maximum of the estimated probability density function.
4 Results and Evaluation
To illustrate what kind of distributions our models [^12] For simplicity, we subsequently use the term “model” in this context to refer to the respective ensemble specification used to generate our global-only, local-only, or global-local predictions. predict, Figure 2 compares observed fatalities during the first test window with our global model predictions for four example grid cells representing typical patterns: A high-intensity case (a, b) and a low-intensity case (c, d) we capture fairly well, and two “misses”: a “false negative” case (e) where our predictions showed only a minuscule chance for violence but in fact a significant number of fatalities occurred, and a “false positive” case (f) where our predictions saw a small but noticeable chance for violence usually seen in low intensity cases, but no violence occurred.
As becomes immediately apparent, our predictions are heavily right-skewed with long tails and a large number of zeroes (a, c), indicating mostly low probabilities in the classification step. Notably, this zero-inflatedness mirrors the distribution of observed fatalities across pgms. The predictions capture general conflict intensity well in most cases, but our hurdle approach only leads to maximum a posteriori (MAP) estimates (see 4.1) above zero for the highest-intensity predictions (a, b). Instead, higher conflict intensity is represented through a greater likelihood of non-zero fatalities. In line with other prediction systems (Mueller and Rauh, 2018; Bazzi et al., 2022), our models are better at predicting the locations of fatalities than their timing, with variance in the predictive distributions over time rarely matching observed fatalities (b, d). However, with the distributional approach, the spikes of violence observed in real-world data can be treated simply as lower likelihood events, but within the range of possible values.
Typical misses include “false negative” cases, where significant violence occurred despite our sample containing more than 99% zeroes, such as in the Cameroon example, located close to the Nigerian border (e). However, our models partially account for this by always assuming a small base probability of violence with the samples never being all-zero, which is the result of our hurdle combination approach and a non-zero minimum probability from the classification step. In addition, we see a number of “false positives” (f), where a sizable share of our samples is non-zero but this predicted risk does not materialize. These predictions generally are part of a spatial cluster of higher-likelihood predictions with fatalities observed only in some of its grid cells. The model thus assumes elevated risk in cells near active conflict and predictions are similar to those where low-intensity violence did occur (d), without being able to reliably identify exact locations.